A concrete monument is a regular pentagonal pyramid with a base edge of
40 cm and one of its lateral edges measure 50 cm.

10. How much concrete is needed to build the monument?
1
Expert's answer
2020-11-19T16:54:34-0500

"SABCDE" is a regular pentagonal pyramid.

"ABCDE" is regular pentagon. "O" is the centre of "ABCDE"

Triangle "BOA"

"AB=40\ cm, \angle AOB=\dfrac{360\degree}{5}=72\degree"

"OK\perp AB, AK=\dfrac{1}{2}AB=20\ cm"


"OA=R=\dfrac{AK}{\sin(\dfrac{\angle AOB}{2})}=\dfrac{20}{\sin(\dfrac{72\degree}{2})}cm"

"\sin36\degree=\dfrac{\sqrt{10-2\sqrt{5}}}{4}"


"OA=\dfrac{80}{\sqrt{10-2\sqrt{5}}}\ cm"

8. Right triangle "SAO"


"H=\sqrt{SA^2-OA^2}"

Given "SA=50\ cm"


"H=\sqrt{(50)^2-(\dfrac{80}{\sqrt{10-2\sqrt{5}}})^2}\ cm="

"=\dfrac{10\sqrt{186-50\sqrt{5}}}{\sqrt{10-2\sqrt{5}}}=10\sqrt{\dfrac{93-25\sqrt{5}}{5-\sqrt{5}}}\approx"

"\approx36.636\ cm"

"S_{ABCDE}=5(\dfrac{1}{2})(OK)(AB)=\dfrac{5}{2}(AK\tan36\degree)(AB)="

"=\dfrac{5}{4}(AB)^2\tan36\degree"

"\sin36\degree=\dfrac{\sqrt{10-2\sqrt{5}}}{4}"

"\cos^2 36\degree=1-\sin^236\degree=1-\dfrac{10-2\sqrt{5}}{16}=\dfrac{6+2\sqrt{5}}{16}"

"\tan^236\degree=\dfrac{10-2\sqrt{5}}{6+2\sqrt{5}}"


"\tan36\degree=\sqrt{\dfrac{5-\sqrt{5}}{3+\sqrt{5}}}"

"S_{ABCDE}=\dfrac{5}{4}(40cm)^2\sqrt{\dfrac{5-\sqrt{5}}{3+\sqrt{5}}}"

"V=\dfrac{1}{3}S_{ABCDE}H="

"=\dfrac{1}{3}(2000)\sqrt{\dfrac{5-\sqrt{5}}{3+\sqrt{5}}}(10\sqrt{\dfrac{93-25\sqrt{5}}{5-\sqrt{5}}})cm^3="

"=\dfrac{20000}{3}\sqrt{\dfrac{93-25\sqrt{5}}{3+\sqrt{5}}}cm^3\approx"

"\approx17745.29cm^3\approx1.8m^3"